PRELIMINARY MEASUREMENTS MADE WITH THE PARIS-IR INSTRUMENT DURING THE ACE ARC-TIC VALIDATION CAMPAIGN

\underline{KEEYOON SUNG}, KALEY A. WALKER, AND PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

The Portable Atmospheric Research Interferometric Spectrometer (PARIS-IR) participated in a field campaign at the Arctic Stratospheric Ozone Observatory at Eureka, Nunavut, Canada (80 N, 86 W) in Winter 2004. The goal of this campaign was to provide measurements to use for validating the instruments on board the Canadian scientific satellite SCISAT-1. The PARIS-IR instrument made solar absorption measurements during the three weeks in February and March following the return of sunlight. This paper will present the preliminary results and analysis from this campaign.